OpenStudy (anonymous):

the population of a local species of dragonfly can be found using an infinite geometric series where a1 = 42 and the common ratio is 3/4. write the sum in sigma notation and calculate the sum that will be the upper limit of this population

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

\[A. \sum_{i=1}^{\infty}(3/4)42^{i-1},\] the sum is divergent \[B. \sum_{i=1}^{\infty}(3/4)42^{i-1}\] the sum is 168 \[C. \sum_{i=1}^{\infty}42(3/4)^{i-1}\] the sum is divergent \[D. \sum_{i=1}^{\infty}42(3/4)^{i-1}\] the sum is 168

OpenStudy (anonymous):

@amorfide

OpenStudy (amorfide):

you know the term a which is 42 you know the ratio r wh ich is 3/4 you know the geometric series to infinity is sum from 1 to infinite and the equation is \[\sum_{1}^{\infty} ar^{k}=\frac{ a }{ 1-r}\]

OpenStudy (amorfide):

I probably should have put k=1 on the bottom of that sigma

OpenStudy (anonymous):

i understand how you got the equation, i just dont know how to solve it :(

OpenStudy (amorfide):

well you would substitute your values a is the first term r is the ratio

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@thomaster

OpenStudy (anonymous):

the answer was D. gee, thanks for the help guys (he said sarcastically)